OpenStudy (anonymous):

The double integral ∫ from 0 to 1 ∫ from x to 1 cos(y^2) dydx ? I have tried doing dx then dy but end up with the same problem of having cos(y^2). Any ideas on what I need to do?

OpenStudy (anonymous):

u substitution i guess.. did u just try it...?

OpenStudy (anonymous):

\[\int\limits_{0}^{1} \cos(y^2)\]

OpenStudy (anonymous):

did i write it correctly?

OpenStudy (anonymous):

Yes

ganeshie8 (ganeshie8):

\[\large \int \limits_0^1 \int \limits_0^x \cos(y^2) ~dy dx\]

ganeshie8 (ganeshie8):

its a double integral, right ?

OpenStudy (anonymous):

sorry i think @ganeshie8 is right.. i forgot that it is double integration....

OpenStudy (anonymous):

oops jnx ganes the x and 0 need to switch

OpenStudy (anonymous):

and it needs to be a one, hold on I need to write it. I was trying to do substitution and am getting confused.

ganeshie8 (ganeshie8):

\[\large \int \limits_0^1 \int \limits_x^1 \cos(y^2) ~dy dx\] like this ?

OpenStudy (anonymous):

\[\int\limits_{0}^{1} \int\limits_{x}^{1} \cos(y ^{2}) dydx\]

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

Substitution tricks will not work

ganeshie8 (ganeshie8):

try to work \(\large dxdy\) instead

ganeshie8 (ganeshie8):

change the order of integration ^

OpenStudy (anonymous):

I did but you end up with the same problem

OpenStudy (anonymous):

It was the first thing I tried

ganeshie8 (ganeshie8):

\[\large \int \limits_0^1 \int \limits_x^1 \cos(y^2) ~dy dx = \large \int \limits_{\cdots}^{\cdots} \int \limits_{\cdots}^{\cdots} \cos(y^2) ~\color{red}{dxdy}\]

OpenStudy (anonymous):

\[\int\limits_{x}^{1} cosy ^{2} dx\]

ganeshie8 (ganeshie8):

\[\large \int \limits_0^1 \int \limits_x^1 \cos(y^2) ~dy dx = \large \int \limits_0^1 \int \limits_0^y \cos(y^2) ~\color{red}{dxdy}\]

ganeshie8 (ganeshie8):

Notice that cos(y^2) is just a constant when you work dx - so you can push it out of integral

ganeshie8 (ganeshie8):

\[\large \int \limits_0^1 \int \limits_x^1 \cos(y^2) ~dy dx = \large \int \limits_0^1 \int \limits_0^y \cos(y^2) ~\color{red}{dxdy}\] \[\large = \large \int \limits_0^1 \cos(y^2) \int \limits_0^y ~\color{red}{dxdy}\]

OpenStudy (anonymous):

but when you substitute 1 and zero in are you not left with cosy^2? or do you not switch the \[\int\limits_{.}^{.}\]

ganeshie8 (ganeshie8):

bounds also change when you change the order of integration

OpenStudy (anonymous):

Ok so I need to find the new bounds?

ganeshie8 (ganeshie8):

Exactly ! but i changed the bounds already, see if that looks okay above ^

OpenStudy (anonymous):

Looks good, thank you for the help!

ganeshie8 (ganeshie8):

np :) evaluating it should be trivial... good luck !